# PACS Collaboration

### v1

@@ II.2.1 Induced pseudoscalar form factor $\widetilde{F}_{P}(q^{2})$
In our previous works Ishikawa:2018rew ; Shintani:2018ozy ; Ishikawa:2021eut ; Tsuji:2023llh , the correlation functions involving the $A_{4}$ current are not taken into account for the calculation of the $\widetilde{F}_{A}(q^{2})$ and $\widetilde{F}_{P}(q^{2})$ form factors. This is simply because, to the best of our knowledge, the $A_{4}$ correlator was found to be statistically very noisy in Ref. Ishikawa:2018rew , where the time-reversal averaging was performed using both forward and backward propagation in time for all three-point correlation functions. However, as pointed out for the first time in Ref. Bali:2018qus , the correlator ratio ${\cal R}^{5z}_{A_{4}}(t,{\bm{q}})$ does not show a plateau, but rather a peculiar behavior that depends linearly on the current insertion time $t$ with a steep negative slope under their kinematic setup. When no time-reversal averaging is applied in our data as shown in Fig. 2, an almost linear $t$ dependence is indeed confirmed, giving the same slope, although the direction is reversed according to the respective kinematics.
@@
Figure 2: Results of the ratio $\mathcal{R}^{5z}_{A_{4}}(t;\bm{q})/(iq_{3}K^{-1})$ calculated from PACS10/L160 with $t_{\mathrm{sep}}/a=13$ (diamonds), $16$ (squares) and $19$ (circles) for seven non-zero momentum transfers (labeled from Q1 to Q7) as functions of the current insertion time slice $t$. The nucleon matrix element of $A_{4}$ exposes the presence of $\pi N$-state contribution described by $\Delta_{-}(t)$ defined in Eq. (24).
@@
As discussed in Refs. Bar:2018xyi ; Bar:2019gfx , such peculiar time dependence is understood as the leading contribution from the $\pi N$ state in ${\cal R}^{5z}_{A_{4}}(t,{\bm{q}})$, arising in the tree diagram of the baryon chiral perturbation theory (ChPT). Importantly, the momentum $\bm{q}$ injected by the axial-vector current is entirely inherited by the pion state, since the pion in such $\pi N$ state remains in the on-mass shell. As illustrated in Fig. 3, the kinematics of the leading correction to the ground-state contribution $N(-\bm{q})\rightarrow N(\bm{0})$ (depicted in Fig. 3(A)) is thus constrained to two special cases Meyer:2018twz ; RQCD:2019jai :
  • $\pi(-\bm{q})N(\bm{0})\rightarrow N(\bm{0})$ (depicted in Fig. 3(B))
  • $N(-\bm{q})\rightarrow\pi(\bm{q})N(-\bm{q})$ (depicted in Fig. 3(C))
@@
where the momentum ${\bm{q}}$ is carried by the axial-vector current operator. It is important to note that neither case is always the lowest energy of the possible $\pi N$ states as excited states. However, in Ref. Jang:2019vkm , a detailed analysis was carried out to identify the main contributions to excited state contamination in ${\cal R}^{5z}_{A_{4}}(t,{\bm{q}})$ using multi-state fits of the two-point function and the $A_{4}$ correlator ${\cal C}^{5z}_{A_{4}}(t,{\bm{q}})$, and indeed the particular $\pi N$ state described in Fig. 3 was confirmed to be a major contribution as the leading $\pi N$ contribution in ${\cal R}^{5z}_{A_{4}}(t,{\bm{q}})$.
@@
Similar contributions from the same $\pi N$-states may appear even in ${\cal R}^{5z}_{A_{i}}(t,{\bm{q}})$. Indeed, Ref. RQCD:2019jai proved it by providing a more general discussion at tree level in baryon ChPT compared to Refs. Bar:2018xyi ; Bar:2019gfx , and showed that the relative signs of the two $\pi N$ contributions are reversed between the $A_{4}$ and $A_{i}$ correlators.
@@
As described earlier at the end of Sec. II.1, in our previous works Ishikawa:2018rew ; Shintani:2018ozy ; Ishikawa:2021eut ; Tsuji:2023llh , the $F_{P}$ form factor obtained from ¯$\overline{\cal R}^{5z}_{A_{i}}(t,{\bm{q}})$ was indeed significantly affected by the excited-state contamination, though no such effect was observed for the $F_{A}$ form factor. Recent studies solving the generalized eigenvalue problem including $\pi N$ operators also show that the $\pi N$ contributions are strong in $\widetilde{F}_{P}(q^{2})$ and $\widetilde{G}_{P}(q^{2})$, but not in $\widetilde{F}_{A}(q^{2})$ Barca:2022uhi ; Alexandrou:2024tin . Therefore, we assume that the contributions from the leading $\pi N$ state for ${\cal R}^{5z}_{A_{\mu}}(t,{\bm{q}})$ can be described as follows
¯$\overline{\cal R}^{5z}_{A_{i}}(t,{\bm{q}})$ $\equiv{\cal R}^{5z}_{A_{i}}(t,{\bm{q}})-\delta_{i3}{\cal R}^{5z}_{A_{3}}(t,{\bm{q}}_{0})$ (20)
@@
For the case when the axial-vector current operator carries the momentum ${\bm{q}}$, the $\pi N$ contribution can be expressed by the following form with $t$-independent coefficients $B(\bm{q})$ and $C(\bm{q})$
@@
which offers us to separate the $\pi N$ contribution $\Delta_{\pm}(t,t_{\mathrm{sep}};{\bm{q}})$ from the $\widetilde{F}_{P}$ form factor using two types of the time-derivative of the correlator ratio, $\partial_{4}\overline{\cal R}^{5z}_{A_{i}}(t,{\bm{q}})$ and $\partial_{4}{\cal R}^{5z}_{A_{4}}(t,{\bm{q}})$  [^3] The time derivative $\partial_{4}$ is defined by $\partial_{4}f(t)=\frac{1}{2}\left(f(t+1)-f(t-1)\right)$ so that Eq. (25) is valid up to ${\cal O}(a^{2})$.. Hereafter the nucleon energy $E_{N}(\bm{q})$ and the pion energy $E_{\pi}(\bm{q})$ are simply abbreviated by shorthand notations $E_{N}$ and $E_{\pi}$, respectively.
@@
with $\Delta E_{N}\equiv E_{N}-M_{N}$ and $K=\sqrt{2E_{N}(E_{N}+M_{N})}$. The first term corresponds to $\widetilde{F}_{P}^{\mathrm{std}}(q^{2})$ in the standard ratio method. The leading $\pi N$ contributions represented in terms of $\Delta_{+}(t,t_{\mathrm{sep}};{\bm{q}})$ and $\Delta_{-}(t,t_{\mathrm{sep}};{\bm{q}})$ can be completely eliminated by adding the second term in Eq. (26). For the ground state contribution, namely $\widetilde{F}_{P}(q^{2})$, Eq. (26) is just a harmless linear combination exploiting the redundancy in the determination of $\widetilde{F}_{P}(q^{2})$ from both of $C_{A_{4}}^{5z}(t;\bm{q})$ and ${C}_{A_{i}}^{5z}(t;\bm{q})$. Furthermore, from Eqs. (21) and (23), it is clear that the terms in the parenthesis in Eq. (26) do not have the ground state contribution. Therefore, if Eq. (26) successfully shows good plateau behavior, independent of the choice of $t_{\mathrm{sep}}$, it guarantees that the ground state contribution can be read accurately without excited-state contamination.
@@ II.2.2 Pseudoscalar form factor $\widetilde{G}_{P}(q^{2})$
where $\Delta_{P}(t,t_{\mathrm{sep}};{\bm{q}})$ encodes the leading $\pi N$ state contributions that cause a residual $t$ dependence in ${\cal R}_{P}^{5z}(t;\bm{q})$ as shown in the lower panel of Fig. 1. Unlike in the case of the axial-vector currents, only a single correlator cannot remove the $\pi N$ contribution $\Delta_{P}(t,t_{\mathrm{sep}};{\bm{q}})$.
It is important to recall here that in our previous study that the axial Ward-Takahashi identity is well satisfied in terms of the three-point functions of the nucleon:
@@
where $m_{\rm PCAC}$ corresponds to the bare quark mass which coincides with the value determined from the pion two-point correlation functions Tsuji:2023llh . It should be emphasized that Eq. (28) is satisfied without isolating the ground-state contribution from the excited-state contributions, as first pointed out in Ref. Bali:2018qus . Thus, Eq. (28) leads to the following PCAC relation for the leading $\pi N$ contributions involved in ¯$\overline{\cal R}_{A_{i}}^{5z}(t;\bm{q})$ and ${\cal R}_{P}^{5z}(t;\bm{q})$:
@@
which also offers a simple subtraction method for determining the $\widetilde{G}_{P}(q^{2})$ as below
@@
with the bare low-energy constant $B_{0}=\frac{M_{\pi}^{2}}{2m_{\mathrm{PCAC}}}$. The first term corresponds to $\widetilde{G}_{P}^{\mathrm{std}}(q^{2})$ in the standard method. See Appendix C for details on the derivation of Eq. (29).
@@ III Numerical results
In the calculation of nucleon two-point and three-point functions, the all-mode-averaging (AMA) technique Blum:2012uh ; Shintani:2014vja ; vonHippel:2016wid is used to significantly reduce the computational cost of multiple measurements and to achieve a much higher statistical accuracy. We compute the combination of the correlation function with the quark propagator in the high-precision and low-precision calculations. For the low-precision calculations, the position of the source operator are changed with respect to the translational symmetry $G$ on the lattice and the temporal direction of the configuration is also changed using the hypercubic symmetry of our lattice setup ($L=T$). If the number of the low-precision calculations is given by $N_{G}$, the total number of measurements (denoted by $N_{\mathrm{meas}}=N_{G}\times N_{\mathrm{conf}}$) is $O(10^{4}\mbox{-}10^{5})$ for the $64^{4}$ lattice or $O(10^{3}\mbox{-}10^{4})$ for the $128^{4}$ and $160^{4}$ lattices. See Refs. Tsuji:2022ric ; Tsuji:2023llh for further details.
@@
Table 3: Summary of results for basic physical quantities in 2+1 flavor PACS10 ensembles with two different lattice spacings: the $\beta$ value, the PCAC quark mass ($\hat{m}_{\mathrm{PCAC}}$:2019ofv ; Tsuji:2023llh , the pion decay constant ($F_{\pi}$shikawa:2022ulx and the nucleon mass ($M_{N}$hintani:2018ozy ; Tsuji:2023llh in lattice units, and the renormalization constant for the axial-vector current determined by the SF scheme ($Z_{A}^{\mathrm{SF}}$:2019ofv ; Tsuji:2023llh .
$\beta$ $a\hat{m}_{\rm PCAC}$  [^1] The bare quark mass $m_{\mathrm{PCAC}}$ that appears in Eq. (28) is defined by $m_{\mathrm{PCAC}}=Z_{A}^{\mathrm{SF}}\hat{m}_{\mathrm{PCAC}}$. $aF_{\pi}$ 222 We use a traditional convention as $F_{\pi}=f_{\pi}/\sqrt{2}\sim 93$ MeV, while $f_{\pi}$ is quoted in Ref. Ishikawa:2022ulx . $aM_{N}$ $Z_{A}^{\mathrm{SF}}$
@@ III.1 Induced pseudoscalar form factor ${F}_{P}(q^{2})$
The $\widetilde{F}_{P}$ form factor is extracted from Eq. (26) as a function of the current insertion time $t$. Figure 4 shows the $t$ dependence and $t_{\mathrm{sep}}$ dependence of $F_{P}(q^{2})=Z_{A}^{\mathrm{SF}}\widetilde{F}_{P}(q^{2})$ obtained by the leading $\pi N$ subtraction method for the $160^{4}$ lattice ensemble (PACS10/L160) as typical examples. The leading $\pi N$ subtraction method is really effective in obtaining an asymptotic plateau in all choices of $t_{\mathrm{sep}}/a=\{13,16,19\}$ for all variation of $q^{2}$. Indeed, the good plateaus show that the $t$ dependence is eliminated and the $t_{\mathrm{sep}}$ dependence is not visible either. Furthermore, the plateau values are consistent with the PPD model estimates given in Eq. (4) within the calculated range of $q^{2}$.
@@
Figure 4: Using the leading $\pi N$ subtraction method as defined in Eq. (26), the values of the $F_{P}$ form factor multiplied by $2M_{N}$ calculated from PACS10/L160 with $t_{\mathrm{sep}}/a=13$ (diamonds), $16$ (squares) and $19$ (circles) for seven non-zero momentum transfers (labeled from Q1 to Q7) as functions of the current insertion time slice $t$. The horizontal dashed line together with yellow band in each panel is calculated from the PPD model ($2M_{N}F_{P}^{\rm PPD}(q^{2})$) given in Eq.(4)
@@
The value of $F_{P}(q^{2})$ can be determined by the uncorrelated constant fit in the wide range between the source and sink points. Figure 5 shows the $t_{\mathrm{sep}}$ dependence of the extracted values of $2M_{N}F_{P}(q^{2})$ for the $128^{4}$ lattice ensemble (left) and the $160^{4}$ lattice ensemble (right). The results given with the different choices of $t_{\mathrm{sep}}$ are mutually consistent with each other within the statistical uncertainties for all seven variations of $q^{2}$. There is no significant $t_{\mathrm{sep}}$ dependence at the same level as was already obtained for the $F_{A}$ form factor Shintani:2018ozy ; Tsuji:2023llh . This indicates that the systematic uncertainties stemming from the excited-state contamination are negligible within the present statistical precision and well under control by the leading $\pi N$ subtraction method together with the optimal choice of the smearing parameter for the nucleon interpolating operator.
@@
Figure 6 shows that the $q^{2}$ dependence of the $F_{P}$ form factor multiplied by $2M_{N}$ for the $128^{4}$ lattice ensemble and the $160^{4}$ lattice ensemble. Two experimental results of the muon capture Gorringe:2002xx ; MuCap:2007tkq ; MuCap:2012lei and the pion electro-production Choi:1993vt are marked as blue diamonds and a black filled circle. Both of our results from the coarse ($128^{4}$) and fine ($160^{4}$) lattices are consistent with each other and agree well with two experimental results. Furthermore, the $q^{2}$ dependence of our results is in good agreement with the prediction of the PPD model given in Eq.(4), using the simulated values of $M_{\pi}$, $M_{N}$, and $F_{A}(q^{2})$.
@@
Figure 5: $t_{\mathrm{sep}}$ dependence for the $F_{P}$ form factor with seven lowest momentum transfers. Results are obtained from the coarse $128^{4}$ lattice (left) and fine $160^{4}$ lattice (right).

### v2

@@ II.2.1 Induced pseudoscalar form factor $\widetilde{F}_{P}(q^{2})$
In our previous works Ishikawa:2018rew ; Shintani:2018ozy ; Ishikawa:2021eut ; Tsuji:2023llh , the correlation functions involving the $A_{4}$ current are not taken into account for the calculation of the $\widetilde{F}_{A}(q^{2})$ and $\widetilde{F}_{P}(q^{2})$ form factors. This is simply because, to the best of our knowledge, the $A_{4}$ correlator was found to be statistically very noisy in Ref. Ishikawa:2018rew , where the time-reversal averaging was performed using both forward and backward propagations in time for all three-point correlation functions. However, as pointed out for the first time in Ref. Bali:2018qus , the correlator ratio ${\cal R}^{5z}_{A_{4}}(t,{\bm{q}})$ does not show a plateau, but rather a peculiar behavior that depends linearly on the current insertion time $t$ with a steep negative slope under their kinematic setup. When no time-reversal averaging is applied in our data as shown in Fig. 2, an almost linear $t$ dependence is indeed confirmed, giving the same slope, although the direction is reversed according to the respective kinematics.
@@
Figure 2: Results of the ratio $\mathcal{R}^{5z}_{A_{4}}(t;\bm{q})/(iq_{3}K^{-1})$ calculated from PACS10/L160 with $t_{\mathrm{sep}}/a=13$ (diamonds), $16$ (squares) and $19$ (circles) for seven nonzero momentum transfers (labeled from Q1 to Q7) as functions of the current insertion time slice $t$. The nucleon matrix element of $A_{4}$ exposes the presence of $\pi N$-state contribution described by $\Delta_{-}(t)$ defined in Eq. (24).
@@
As discussed in Refs. Bar:2018xyi ; Bar:2019gfx , such peculiar time dependence is understood as the leading contribution from the $\pi N$ state in ${\cal R}^{5z}_{A_{4}}(t,{\bm{q}})$, arising in the tree diagram of the baryon chiral perturbation theory (ChPT). Importantly, the momentum $\bm{q}$ injected by the axial-vector current is entirely inherited by the pion state, since the pion in such $\pi N$ state remains in the on-mass shell. As illustrated in Fig. 3, the kinematics of the leading correction to the ground-state contribution $N(-\bm{q})\rightarrow N(\bm{0})$ [depicted in Fig. 3(A)] is thus constrained to two special cases Meyer:2018twz ; RQCD:2019jai :
  • $\pi(-\bm{q})N(\bm{0})\rightarrow N(\bm{0})$ [depicted in Fig. 3(B)]
  • $N(-\bm{q})\rightarrow\pi(\bm{q})N(-\bm{q})$ [depicted in Fig. 3(C)]
@@
where the momentum $\bm{q}$} is carried by the axial-vector current operator. It is important to note that neither case is always the lowest energy of the possible $\pi N$ states as excited states. However, in Ref. Jang:2019vkm , a detailed analysis was carried out to identify the main contributions to excited-state contamination in ${\cal R}^{5z}_{A_{4}}(t,{\bm{q}})$ using multi-state fits of the two-point function and the $A_{4}$ correlator ${\cal C}^{5z}_{A_{4}}(t,{\bm{q}})$, and indeed the particular $\pi N$ state described in Fig. 3 was confirmed to be a major contribution as the leading $\pi N$ contribution in ${\cal R}^{5z}_{A_{4}}(t,{\bm{q}})$.
@@
Similar contributions from the same $\pi N$ states may appear even in ${\cal R}^{5z}_{A_{i}}(t,{\bm{q}})$. Indeed, Ref. RQCD:2019jai proved it by providing a more general discussion at tree level in baryon ChPT compared to Refs. Bar:2018xyi ; Bar:2019gfx , and showed that the relative signs of the two $\pi N$ contributions are reversed between the $A_{4}$ and $A_{i}$ correlators.
@@
As described earlier at the end of Sec. II.1, in our previous works Ishikawa:2018rew ; Shintani:2018ozy ; Ishikawa:2021eut ; Tsuji:2023llh , the $F_{P}$ form factor obtained from $\overline{\cal R}^{5z}_{A_{i}}(t,{\bm{q}})$ was indeed significantly affected by the excited-state contamination, though no such effect was observed for the $F_{A}$ form factor. Recent studies solving the generalized eigenvalue problem including $\pi N$ operators also show that the $\pi N$ contributions are strong in $\widetilde{F}_{P}(q^{2})$ and $\widetilde{G}_{P}(q^{2})$, but not in $\widetilde{F}_{A}(q^{2})$ Barca:2022uhi ; Alexandrou:2024tin . Therefore, we assume that the contributions from the leading $\pi N$ state for ${\cal R}^{5z}_{A_{\mu}}(t,{\bm{q}})$ can be described as follows
$\overline{\cal R}^{5z}_{A_{i}}(t,{\bm{q}})$ $\equiv{\cal R}^{5z}_{A_{i}}(t,{\bm{q}})-\delta_{i3}{\cal R}^{5z}_{A_{3}}(t,{\bm{q}}_{0})$ (20)
@@
For the case when the axial-vector current operator carries the momentum $\bm{q}$}, the $\pi N$ contribution can be expressed by the following form with $t$-independent coefficients $B(\bm{q})$ and $C(\bm{q})$:
@@
which offers us to separate the $\pi N$ contribution $\Delta_{\pm}(t,t_{\mathrm{sep}};{\bm{q}})$ from the $\widetilde{F}_{P}$ form factor using two types of the time derivative of the correlator ratio, $\partial_{4}\overline{\cal R}^{5z}_{A_{i}}(t,{\bm{q}})$ and $\partial_{4}{\cal R}^{5z}_{A_{4}}(t,{\bm{q}})$ [^3] The time derivative $\partial_{4}$ is defined by $\partial_{4}f(t)=\frac{1}{2}\left(f(t+1)-f(t-1)\right)$ so that Eq. (25) is valid up to ${\cal O}(a^{2})$. Hereafter the nucleon energy $E_{N}(\bm{q})$ and the pion energy $E_{\pi}(\bm{q})$ are simply abbreviated by shorthand notations $E_{N}$ and $E_{\pi}$, respectively.
@@
with $\Delta E_{N}\equiv E_{N}-M_{N}$ and $K=\sqrt{2E_{N}(E_{N}+M_{N})}$. The first term corresponds to $\widetilde{F}_{P}^{\mathrm{std}}(q^{2})$ in the standard ratio method. The leading $\pi N$ contributions represented in terms of $\Delta_{+}(t,t_{\mathrm{sep}};{\bm{q}})$ and $\Delta_{-}(t,t_{\mathrm{sep}};{\bm{q}})$ can be completely eliminated by adding the second term in Eq. (26). For the ground-state contribution, namely $\widetilde{F}_{P}(q^{2})$, Eq. (26) is just a harmless linear combination exploiting the redundancy in the determination of $\widetilde{F}_{P}(q^{2})$ from both of $C_{A_{4}}^{5z}(t;\bm{q})$ and ${C}_{A_{i}}^{5z}(t;\bm{q})$. Furthermore, from Eqs. (21) and (23), it is clear that the terms in the parentheses in Eq. (26) do not have the ground-state contribution. Therefore, if Eq. (26) successfully shows good plateau behavior, independent of the choice of $t_{\mathrm{sep}}$, it guarantees that the ground-state contribution can be read accurately without excited-state contamination.
@@ II.2.2 Pseudoscalar form factor $\widetilde{G}_{P}(q^{2})$
where $\Delta_{P}(t,t_{\mathrm{sep}};{\bm{q}})$ encodes the leading $\pi N$-state contributions that cause a residual $t$ dependence in ${\cal R}_{P}^{5z}(t;\bm{q})$ as shown in the lower panel of Fig. 1. Unlike in the case of the axial-vector currents, only a single correlator cannot remove the $\pi N$ contribution $\Delta_{P}(t,t_{\mathrm{sep}};{\bm{q}})$.
It is important to recall here that in our previous study that the axial Ward-Takahashi identity is well satisfied in terms of the three-point functions of the nucleon, as follows:
@@
where $m_{\rm PCAC}$ corresponds to the bare quark mass which coincides with the value determined from the pion two-point correlation functions Tsuji:2023llh . It should be emphasized that Eq. (28) is satisfied without isolating the ground-state contribution from the excited-state contributions, as first pointed out in Ref. Bali:2018qus . Thus, Eq. (28) leads to the following PCAC relation for the leading $\pi N$ contributions involved in $\overline{\cal R}_{A_{i}}^{5z}(t;\bm{q})$ and ${\cal R}_{P}^{5z}(t;\bm{q})$:
@@
which also offers a simple subtraction method for determining the $\widetilde{G}_{P}(q^{2})$, as follows:
@@
with the bare low-energy constant $B_{0}=\frac{M_{\pi}^{2}}{2m_{\mathrm{PCAC}}}$. The first term corresponds to $\widetilde{G}_{P}^{\mathrm{std}}(q^{2})$ in the standard ratio method. See Appendix C for details on the derivation of Eq. (29).
@@ III Numerical results
In the calculation of nucleon two-point and three-point functions, the all-mode-averaging (AMA) technique Blum:2012uh ; Shintani:2014vja ; vonHippel:2016wid is used to significantly reduce the computational cost of multiple measurements and to achieve a much higher statistical accuracy. We compute the combination of the correlation function with the quark propagator in the high-precision and low-precision calculations. For the low-precision calculations, the position of the source operator is changed with respect to the translational symmetry $G$ on the lattice and the temporal direction of the configuration is also changed using the hypercubic symmetry of our lattice setup ($L=T$). If the number of the low-precision calculations is given by $N_{G}$, the total number of measurements (denoted by $N_{\mathrm{meas}}=N_{G}\times N_{\mathrm{conf}}$) is $O(10^{4}\mbox{-}10^{5})$ for the $64^{4}$ lattice or $O(10^{3}\mbox{-}10^{4})$ for the $128^{4}$ and $160^{4}$ lattices. See Refs. Tsuji:2022ric ; Tsuji:2023llh for further details.
@@
Table 3: Summary of results for basic physical quantities in 2+1 flavor PACS10 ensembles with two different lattice spacings: the $\beta$ value, the PCAC quark mass ($\hat{m}_{\mathrm{PCAC}}$PACS:2019ofv ; Tsuji:2023llh , the pion decay constant ($F_{\pi}$Ishikawa:2022ulx and the nucleon mass ($M_{N}$Shintani:2018ozy ; Tsuji:2023llh in lattice units, and the renormalization constant for the axial-vector current determined by the SF scheme ($Z_{A}^{\mathrm{SF}}$PACS:2019ofv ; Tsuji:2023llh .
$\beta$ $a\hat{m}_{\rm PCAC}$  [^1] The bare quark mass $m_{\mathrm{PCAC}}$ that appears in Eq. (28) is defined by $m_{\mathrm{PCAC}}=Z_{A}^{\mathrm{SF}}\hat{m}_{\mathrm{PCAC}}$. $aF_{\pi}$  [^2] We use a traditional convention as $F_{\pi}=f_{\pi}/\sqrt{2}\sim 93$ MeV, while $f_{\pi}$ is quoted in Ref. Ishikawa:2022ulx . $aM_{N}$ $Z_{A}^{\mathrm{SF}}$
@@ III.1 Induced pseudoscalar form factor ${F}_{P}(q^{2})$
The $\widetilde{F}_{P}$ form factor is extracted from Eq. (26) as a function of the current insertion time $t$. Figure 4 shows the $t$ dependence and $t_{\mathrm{sep}}$ dependence of $F_{P}(q^{2})=Z_{A}^{\mathrm{SF}}\widetilde{F}_{P}(q^{2})$ obtained by the leading $\pi N$ subtraction method for the $160^{4}$ lattice ensemble (PACS10/L160) as typical examples. The leading $\pi N$ subtraction method is really effective in obtaining an asymptotic plateau in all choices of $t_{\mathrm{sep}}/a=\{13,16,19\}$ for all variations of $q^{2}$. Indeed, the good plateaus show that the $t$ dependence is eliminated and the $t_{\mathrm{sep}}$ dependence is not visible either. Furthermore, the plateau values are consistent with the PPD model estimates given in Eq. (4) within the calculated range of $q^{2}$.
@@
Figure 4: Using the leading $\pi N$ subtraction method as defined in Eq. (26), the values of the $F_{P}$ form factor multiplied by $2M_{N}$ calculated from PACS10/L160 with $t_{\mathrm{sep}}/a=13$ (diamonds), $16$ (squares) and $19$ (circles) are shown for seven nonzero momentum transfers (labeled from Q1 to Q7) as functions of the current insertion time slice $t$. The horizontal dashed line together with yellow band in each panel is calculated from the PPD model ($2M_{N}F_{P}^{\rm PPD}(q^{2})$) given in Eq.(4)
@@
The value of $F_{P}(q^{2})$ can be determined by the uncorrelated constant fit in the wide range between the source and sink points. Figure 5 shows the $t_{\mathrm{sep}}$ dependence of the extracted values of $2M_{N}F_{P}(q^{2})$ for the $128^{4}$ lattice ensemble (left) and the $160^{4}$ lattice ensemble (right). The results given with the different choices of $t_{\mathrm{sep}}$ are mutually consistent with each other within the statistical uncertainties for all seven variations of $q^{2}$. There is no significant $t_{\mathrm{sep}}$ dependence at the same level as was already obtained for the $F_{A}$ form factor Shintani:2018ozy ; Tsuji:2023llh . We also compare them to the results obtained using the summation method Maiani:1987by to assess residual excited-state contamination after removing the leading $\pi N$ contribution. The results obtained by the two methods are consistent with each other within the statistical uncertainties. This indicates that the systematic uncertainties stemming from the excited-state contamination are negligible within the present statistical precision and well under control by the leading $\pi N$ subtraction method together with the optimal choice of the smearing parameter for the nucleon interpolating operator. See, Appendix G for details of the summation analysis.
@@
Figure 6 shows the $q^{2}$ dependence of the $F_{P}$ form factor multiplied by $2M_{N}$ for the $128^{4}$ lattice ensemble and the $160^{4}$ lattice ensemble. Two experimental results of the muon capture Gorringe:2002xx ; MuCap:2007tkq ; MuCap:2012lei and the pion electro-production Choi:1993vt are marked as blue diamonds and a black filled circle. Both of our results from the coarse ($128^{4}$) and fine ($160^{4}$) lattices are consistent with each other and agree well with two experimental results. Furthermore, the $q^{2}$ dependence of our results is in good agreement with the prediction of the PPD model given in Eq. (4), using the simulated values of $M_{\pi}$, $M_{N}$, and $F_{A}(q^{2})$. In addition, our subtraction method can get more accurate results of $F_{P}(q^{2})$, than those from the multi-state fit approach that was used in other groups Gupta:2024qip . See Appendix D for more details.
@@
Figure 5: $t_{\mathrm{sep}}$ dependence on the $F_{P}$ form factor is shown for seven lowest momentum transfers, along with a comparison of the summation method. Results are obtained from the coarse $128^{4}$ lattice (left) and fine $160^{4}$ lattice (right).